OpenStudy (anonymous):

1/8 of tony's books are mytery books.he has 3 mystery books.how many books does tony have in all?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

1/8 of x = 3 (" of " means multiply) (1/8)x = 3 now you have to divide 3 by 1/8 to get x by itself) 3 divided by 1/8 = 3 x 8/1 = 24 < -- answer

OpenStudy (sepeario):

You might get some more help if you posted your answer in the Mathematics section.. @ezekielrivas